OpenStudy (queenv):

Y-1= - 3 — — X+3 4 4(y-3)=-3(x+4) = 4y-12 = -3x + 1 = 4y -12 + 3 = -3x +3 +1 = 4y -9 = 1 = 4y – 9 -1=1-1 =4y + 10 = 4y + 10 ..4…….4 Y= 10 = 5 — — 4 2 Is this correct? This is Algebra II Ratios and proportions.

OpenStudy (shamil98):

\[\large \frac{ y-1 }{ x+3 } = -\frac{ 3 }{ 4 }\] is this your problem?..

OpenStudy (shamil98):

Hello?.. Are you solving for x? for y? What are you supposed to do?

OpenStudy (queenv):

both and yes

OpenStudy (queenv):

solving for y

OpenStudy (shamil98):

Okay.

OpenStudy (shamil98):

Cross multiplying. 4(y-1) = -3(x+3) distribute. 4y - 4 = -3x - 9 4y = -3x - 5 4y = -3x - 5 y = (-3/4)x - 5/4

OpenStudy (queenv):

okay I am not understanding the answer. How did you get the answer?

OpenStudy (shamil98):

I divided everything by 4. to get y by itself.

OpenStudy (queenv):

I see, I did everything backwards.

OpenStudy (queenv):

Thanks

OpenStudy (queenv):

How do I give you a medal?

OpenStudy (queenv):

I did it